Opunui Corporation has two manufacturing departments --Molding and Finishing. The company used the following data at the begiMultiple Choice 0 $8,800 $9,000 $30,986 $13,186
Answer #1

Predetermined overhead rate for modeling = 25000/4000 = $6.25 per hour

Predetermined overhead rate for Finishing = 3700/1000 = $3.7 per hour

Job M
Direct material $8800
Direct labor 9000
Fixed manufacturing overhead (6.25*1300 + 3.7*600) 10345
Variable manufacturing overhead (1300*1 + 600*2) 2500
Total manufacturing cost $30645
Answer #2

Predetermined   OH Rate = Total estimated fixed Manufacturing OH / Total estimated machine   hours




Estimated fixed manuf. OH           28,700
Add: Variable OH-Molding             4,000(1*4000)
Add: Variable OH-Fabricating             2,000(2*1000)



Total Overhead cost           34,700
Divided by estimated Total Machine hours             5,000



Predetermined OH Rate                6.94



Manufacturing overhead applied to Job A   and Job M


Job M



Machine hours actually used             1,900            
Multiplied by Plantwide OH   Rate                6.94               



Total Overhead apply to Jobs           13,186          



Total amount of manufacturing cost   assigned to Job M




Direct Material             8,800
Direct Labour             9,000
Overhead applied            13,186



Total amount of manufacturing   cost assigned to Job M           30,986
CORRECT (C)


c:30986


source: managerial accounting
answered by: anonymous